A car initially traveled at a speed of 16 meters per second acceleration uniformly to a speed of 20 meters per second over a distance of 36 meters what is the acceleration


1
Expert's answer
2020-09-30T12:22:15-0400

The acceleration is


"a=\frac{v_f^2-v_i^2}{2d},\\\space\\
a=\frac{20^2-16^2}{2\cdot36}=2\text{ m/s}^2."